Circular No. 7830

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2000fs IN NGC 1218 A. R. Martel, Johns Hopkins University, and collaborators report the

discovery of a supernova (V = 18.7) in NGC 1218 (z = 0.0286), which is host to the radio source 3C 78. The supernova was detected on 2000 Sept. 6 UT in an acquisition image (filter F28X50LP) and spectrum of the galaxy’s core, obtained as part of the Hubble Space Telescope (HST) spectroscopic program GO-8700 with the Space Telescope Imaging Spectrograph (STIS). SN 2000fs is located at α = 3^{h}08^{m}26^{s}.24, δ = +4^{o}06′39′′.9 (equinox 2000.0), which is ≈ 0′′.5 northeast of the galaxy’s nucleus. The new object is superimposed along the northern edge of the optical synchrotron jet of 3C 78. Comparison with HST WFPC2 images taken on 1994 Aug. 17 (red F702W filter) and 1996 Sept. 16 (blue F555W filter) shows no evidence of the supernova. The STIS spectra (range 282.5–998.0 nm) of the supernova appear generally consistent with a type-Ia event, with the Ca II triplet, Mg II, Si II, and Fe II lines detected.

SUPERNOVA 2001gd IN NGC 5033 C. J. Stockdale, Naval Research Laboratory (NRL); M. A. Pérez-

Torres, Istituto di Radioastronomia, Bologna; J. M. Marcaide, University of Valencia; R. A. Sramek, National Radio Astronomy Observatory; K. W. Weiler, NRL; S. D. Van Dyk, Infrared Processing and Analysis Center, California Institute of Technology; N. Panagia, European Space Agency and Space Telescope Science Institute; P. Lundqvist, Stockholm University; D. Pooley, Massachusetts Institute of Technology (MIT); S. Immler, University of Massachusetts, Amherst; and W. Lewin, MIT, report the detection of radio emission from the type-IIb supernova 2001gd (IAUC 7761, 7765) with the Very Large Array. Radio-flux densities on Feb. 8.54 UT of 1.37 ± 0.18 mJy at 22.485 GHz (wavelength 1.3 cm), 4.09 ± 0.21 mJy at 8.435 GHz (wavelength 3.6 cm), 5.05 ± 0.26 mJy at 4.885 GHz (wavelength 6.1 cm), and 0.85 ± 0.13 mJy at 1.465 GHz (wavelength 20.5 cm) were measured in ‘A’ configuration at the following position: α = 13^{h}13^{m}23^{s}.899, δ = +36^{o}38′18′′.14 (equinox J2000; ± 0′′.2 in each coordinate). This is in close agreement with the reported optical position (end figures 23^{s}.89, 17′′.7; IAUC 7761). NGC 5033 is also the host galaxy of SNe 1985L and 1950C, from which no radio emission was detected in these observations (with 1.465 GHz 3σ upper limits of 0.261 mJy to each observation).

2002 February 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green